OpenStudy (anonymous):

Help me please i only want the answer ill fan u and give u a medal TY Admission to the Pie Eating Contest is $1.50 for children and $4 for adults. On a certain day, 2200 people came to watch and $5050 is collected. How many children and adults attended the Contest? Show ALL work. Let x = number of children Let y = number of adults A. Write 2 equations to model this system: B. Solve the first equation for x (get the x alone on 1 side): 5 points C. Use Substitution to solve the equations and determine how many adults and how many children attended

OpenStudy (anonymous):

help

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

What are you working on I gotta know what type of problem to make

OpenStudy (anonymous):

i need to explain the work and find out Admission to the Pie Eating Contest is $1.50 for children and $4 for adults. On a certain day, 2200 people came to watch and $5050 is collected. How many children and adults attended the Contest? Show ALL work. Let x = number of children Let y = number of adults A. Write 2 equations to model this system: B. Solve the first equation for x (get the x alone on 1 side): 5 points C. Use Substitution to solve the equations and determine how many adults and how many children attended

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

your problem can be modeled by the subsequent system: \[\left\{ {\begin{array}{*{20}{c}} {1.5x + 4y = 5050} \\ {x + y = 2200} \end{array}} \right.\] the first equation comes from the total money collected, the second one comes from the total number of person who wants to assist to the performance

OpenStudy (anonymous):

so is that for part a ?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

from the second equation I get: x= 2200-y, now I substitute into the first one so I get: 1.5(2200-y)+4*y=5050 please solve for y

OpenStudy (anonymous):

4(2200-x)+1.5*x=5050 ? is that correct ?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

what is x?

OpenStudy (anonymous):

the amount of children

OpenStudy (anonymous):

now can u help with part C pls and TY

OpenStudy (michele_laino):

we have to simplify that equation, so we get: 8800-4x+1.5x=5050 please continue

OpenStudy (anonymous):

how did u get 8800 ?

OpenStudy (anonymous):

i dont get it

OpenStudy (michele_laino):

4*2200= 8800

OpenStudy (anonymous):

oh ok i get it TY

OpenStudy (anonymous):

so 3300-1.5x+4y=5500 am i right ?

OpenStudy (michele_laino):

after that simplification, you should get this equation: 2.5 x= 3750

OpenStudy (michele_laino):

more steps: 8800-2.5x=5050

OpenStudy (michele_laino):

2.5 x= 8800-5050

OpenStudy (michele_laino):

so: x=3750/2.5=...?

OpenStudy (michele_laino):

what is: \[\Large x = \frac{{3750}}{{2.5}} = ...?\]

OpenStudy (anonymous):

1500

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now, we have: y=2200-x=2200-1500=...?

OpenStudy (anonymous):

700

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

ty soooooo much can you organize according to the answers to the Parts pls ill right a review and fan and medal you ty again

OpenStudy (michele_laino):

part A: the requested system is: \[\left\{ {\begin{array}{*{20}{c}} {1.5x + 4y = 5050} \\ {x + y = 2200} \end{array}} \right.\]

OpenStudy (michele_laino):

part B: the requested equation is: \[2.5x = 3750\]

OpenStudy (michele_laino):

part C: the requested solution is: \[\left\{ {\begin{array}{*{20}{c}} {x = 1500} \\ {y = 700} \end{array}} \right.\]

OpenStudy (anonymous):

yaaaaaaay TY sooooo sooooooo sooooooo much im happy i found you

OpenStudy (michele_laino):

thanks!!!!! :)

OpenStudy (anonymous):

np